### v4

# Revisiting Optimal Allocations for Binary Responses: Insights from Considering Type-I Error Rate Control

Abstract
This work revisits optimal response-adaptive designs from a type-I error rate perspective, highlighting when and how much these allocations exacerbate type-I error rate inflation — an issue previously undocumented. We explore a range of approaches from the literature that can be applied to reduce type-I error rate inflation. However, we found that all of these approaches fail to give a robust solution to the problem. To address this, we derive two optimal allocation proportions, incorporating the more robust score test (instead of the Wald test) with finite sample estimators (instead of the unknown true values) in the formulation of the optimization problem. One proportion optimizes statistical power and the other minimizes the total number failures in a trial while maintaining a fixed variance level. Through simulations based on an early-phase and a confirmatory trial we provide crucial practical insight into how these new optimal proportion designs can offer substantial patient outcomes advantages while controlling type-I error rate. While we focused on binary outcomes, the framework offers valuable insights that naturally extend to other outcome types, multi-armed trials and alternative measures of interest.
Keywords Neyman allocation; patient benefit; RSHIR allocation; score test; Wald test.
1 Introduction
While adaptive designs hold promise, their practical application in confirmatory trials lags behind their theoretical prominence (e.g., Burnett et al., 2020; Dimairo et al., 2015). Response-adaptive randomization (RAR) designs is an example of this phenomenon, having a history spanning nearly a century (Thompson, 1933) but remaining more a topic of persistent controversy among statisticians (Robertson et al., 2023) than a practically chosen design option. This is partly due to the need to balance trade-offs in terms of patient-benefit considerations, power, and type-I error rate as part of achieving acceptability for their practical application. Often discussions of the use of RAR have focused on the trade-off between patient-benefit metrics and statistical power while the type-I error rate has been ignored in the theoretical literature. From a regulatory perspective, type-I error rate control is an absolute minimal requirement for any trial design, without which concerns about power or patient benefit become tangential.
In this paper, we provide new insights for clinical trial designers by proposing two novel optimal RAR designs that explicitly consider type-I error rate control alongside power and patient benefits. Our focus is two-armed trials with binary endpoints, which are commonly used in clinical research. In a literature review on PubMed, we identified over 2,200 papers that discuss Phase III trials with a binary endpoint published in the last 10 years. We reexamine from a new perspective the paper by Rosenberger et al. (2001), who first proposed an approach based on targeting optimal allocation proportions, and is one of the most cited papers (335) in the RAR literature.
In Section 2, we present two optimal allocation proportions that were presented in Rosenberger et al. (2001) and serve as the foundation of our analysis. The first, Neyman allocation, maximizes the statistical power of a widely used Wald test. The second, RSHIR allocation, an abbreviation of the names of authors of Rosenberger et al. (2001), minimizes the expected number of patient failures during the trial while ensuring the test maintains a predefined variance level. Both allocations are derived for the Wald test. Other have reported issues with its use as an inferential method following response-adaptive trials (Baldi Antognini et al., 2018). Here, we reveal a novel finding: using Neyman or RSHIR allocation further exacerbates type-I error rate inflation to a degree that may surprise readers. To our knowledge, this issue has not been documented in the prior literature.
Section 3 illustrates the type-I error rate performance of approaches that have originally been suggested for related issues, such as non-zero variance estimation. However, none of these approaches lead to satisfactory type-I error rate control. To address these concerns, in Section 4 we explore how to derive optimal proportions for an alternative test statistic and demonstrate how the principles behind the two optimal proportions can be adapted to mitigate statistical issues. In Section 5, we provide practical guidance for using optimal proportions in designing some clinical trial examples, highlighting their potential advantages and limitations. Finally, Section 6 concludes with a discussion of the implications of our findings for both the design and evaluation of RAR-based trials.
Our main contribution is the development of a procedure and test statistic that preserves the type-I error rate for binary response RAR designs. Note that the structure of this paper follows the way in which we arrived at this procedure. We emphasize, that in practice, one should first ensure that type-I error rates are controlled before power and patient-benefit gains are considered.
2 Notation and Framework for Optimal Response Adaptive Randomization
In this and the next two sections, we consider a two-armed binary trial with a fixed sample size $n=n_{0}+n_{1}$, where $n_{0}$ and $n_{1}$ are the sample sizes for the control and treatment arms, respectively. The responses of individual patients are modeled as random variables $Y_{ki}$. Here, $k\in\{0,1\}$ indicates the treatment ($k=1$) or control ($k=0$), and $i\in\{1,\dots,n\}$ indexes the patients. For each arm $k\in\{0,1\}$, the random variables $Y_{k1},...,Y_{kN}$ are assumed to be independent and identically distributed (i.i.d.) as Bernoulli random variables
$Y_{ki}\mathop{\sim}\limits^{\mathrm{i.i.d.}}\text{Bern}(p_{k}),\quad i=1,\dots,n_{k};\quad k=0,1;\quad n_{0}+n_{1}=n,$
where $p_{k}$ represents the success probability of arm $k$, and $q_{k}=1-p_{k}$ is the probability of failure.
Out of the $2\times n$ potential outcomes (corresponding to two treatments and $n$ patients), only $n$ are observed in the trial, since for each patient we can observe the outcome under only one treatment. The treatment assignment for patient $i$ is indicated by a binary variable $A_{i}$, where $A_{i}=1$ if the patient is assigned to the experimental arm and $A_{i}=0$ otherwise. The probability of assigning a patient to treatment $k$ is therefore $P(A_{i}=1)$, and $P(A_{i}=0)=1-P(A_{i}=1)$. Because randomization minimizes bias, reduces confounding, and ensures comparability between treatment groups, we focus here only on designs that satisfy: $P(A_{i}=k)>0$ for all $i=1,\dots,n$ and $k=0,1$. The simplest (non-adaptive) randomization scheme is complete randomization (CR), where $P(A_{i}=k)=1/2$ for all $i$ and $k$ in a two-arm trial. CR attains an equal allocation proportion (on average) and it is often assumed to maximize statistical power (Berger et al., 2021; Friedman et al., 2015) and control the type-I error rate.
In contrast, RAR dynamically adjusts allocation probabilities based on accumulating trial data. We assume patients enter the trial sequentially, with outcomes observed immediately, to explore the full potential of RAR designs. Let $a^{(j)}=\{a_{1},\dots,a_{j}\}$ and $y^{(j)}=\{y_{1},\dots,y_{j}\}$ represent the allocation and response histories after $j$ patients, respectively, with $a^{(0)}$ and $y^{(0)}$ as empty sets. Under RAR, the allocation probability for patient $i$ depends on prior allocations and responses $P(A_{i}=1|a^{(i-1)},y^{(i-1)})$ for $i=2,\dots,n$. For at least the first patient, the probabilities are typically set to $P(A_{1}=1)=1/2$, reflecting initial equipoise. However, prior knowledge from earlier studies can inform alternative initial probabilities.
At the trial’s conclusion, a statistical test is chosen for final inference. The two-sample Wald test is widely used in RAR literature for evaluating optimal designs (Hu and Rosenberger, 2006). This test considers a measure of interest based on the success probabilities, denoted as $f(p_{0},p_{1})$. In this paper, we focus (as does Rosenberger et al. (2001)) on the mean parameter difference $f(p_{0},p_{1})=p_{1}-p_{0}$, though other measures such as the log relative risk or log odds ratio can also be used as suggested by Rosenberger et al. (2001). A complete table of optimal allocation proportions is given by Pin et al. (2024, p. 325). The general null hypothesis is
$H_{0}:f(p_{0},p_{1})=f_{0}\quad\text{versus}\quad H_{1}:f(p_{0},p_{1})\neq f_{0},$
which can be evaluated using the test statistic
$Z=\frac{f(\hat{p}_{0},\hat{p}_{1})-f_{0}}{\sqrt{\widehat{\mathrm{Var}}(f(\hat{p}_{0},\hat{p}_{1}))}},$ (1)
where $\hat{p}_{k}$ are the MLEs of $p_{k}$, and $\widehat{\mathrm{Var}}(f(\hat{p}_{0},\hat{p}_{1}))$ is the estimated variance of the test statistic. For the mean difference $f(p_{0},p_{1})=p_{1}-p_{0}$ and $f_{0}=0$, the test statistic (1) simplifies to
$Z_{1}=\frac{\hat{p}_{1}-\hat{p}_{0}}{\sqrt{\frac{\hat{p}_{0}\hat{q}_{0}}{n_{0}}+\frac{\hat{p}_{1}\hat{q}_{1}}{n_{1}}}}.$
In practice, RAR is typically combined with early stopping and allows for the allocation of patients in groups and delayed outcomes. However, to fully explore the potential of these algorithms, we will use the framework described above. In the next section, we present two optimal sampling proportions for the test statistic $Z_{1}$, as derived by Tschuprow (1923); Neyman (1934); Robbins (1952) and Rosenberger et al. (2001).
2.1 Two Optimal Proportions for the Wald Test
The first allocation we discuss is Neyman allocation (Neyman, 1934; Robbins, 1952). The aim is to fix the asymptotic variance of the the treatment difference estimator (or equivalently the inverse of the Fisher information, as in Jennison and Turnbull’s formulation (2000, p. 327)), given by $s_{T}^{2}=(p_{0}q_{0})/n_{0}+(p_{1}q_{1})/n_{1}$ (Rosenberger et al., 2001, pp. 910), and under that constraint optimize
$\text{min}_{\rho}n_{0}+n_{1}\textbf{ \quad s.t. \quad}s_{T}^{2}\leq C,$
where $(1-\rho)=n_{0}/n$, $\rho=n_{1}/n$ and $C\in\mathbb{R}^{+}$. Deriving with respect to $\rho$ yields
$\rho_{N_{1}}=\frac{\sqrt{p_{1}q_{1}}}{\sqrt{p_{0}q_{0}}+\sqrt{p_{1}q_{1}}},$
Neyman allocation for the Wald type test $Z_{1}$.
The second optimal allocation we introduce is RSHIR allocation (Rosenberger et al., 2001), which minimizes the expected number of failures in a trial with binary responses, subject to a constraint on the variance of the Wald test. This allocation is determined by solving the following optimization problem, ensuring the variance remains below a pre-specified threshold $C$:
$\text{min}_{\rho}\,n_{0}(1-p_{0})+n_{1}(1-p_{1})\quad\text{subject to}\quad s_{T}^{2}\leq C,$
where $n_{0}=(1-\rho)\cdot n$, $n_{1}=\rho\cdot n$, and $C\in\mathbb{R}^{+}$. Solving the optimization problem leads to the RSHIR allocation for the Wald test $Z_{1}$
$\rho_{R_{1}}=\frac{\sqrt{p_{1}}}{\sqrt{p_{0}}+\sqrt{p_{1}}}.$
Notice that to derive the optimal proportions, we utilized the theoretical variance of the test statistic $Z_{1}$ and the true success probabilities. However, these parameters are typically unknown in a real trial setting.
2.2 Targeting Procedures
If the true parameters $p_{0}$ and $p_{1}$ were known, they could be directly substituted into the optimal allocation proportions $\rho_{N_{1}}$ and $\rho_{R_{1}}$, to guide the sampling process throughout the trial. However, since these parameters are typically unknown, they must be estimated.
One common approach is to use MLEs to obtain parameter estimates $\hat{p}_{0}$ and $\hat{p}_{1}$. In the sequential maximum likelihood procedure (SMLE), the parameters are re-estimated sequentially after observing the outcomes of each patient. Before assigning a patient $j\in\{B,\ldots,n-1\}$, the current estimates $\hat{p}_{0}(j)$ and $\hat{p}_{1}(j)$ are plugged into the optimal allocation proportion formula to compute an updated sampling proportion, denoted as $\hat{\rho}_{N_{1}}(j)$ or $\hat{\rho}_{R_{1}}(j)$. This proportion is then used to allocate the next patient $j+1$.
Two more recently developed methods are the doubly adaptive biased coin design (DBCD), introduced by Eisele (1994); Eisele and Woodroofe (1995) and modified by Hu and Zhang (2004); and the efficient randomized-adaptive design (ERADE) introduced by Hu et al. (2009). ERADE is the discretized adaptation of the doubly adaptive biased coin design (DBCD) by Eisele (1994) aiming to decrease the variance of the design. Given a parameter $\alpha\in(0,1)$, the probability of assigning patient $j+1$ to treatment 1 is determined as follows:
$\rho(n_{1}(j),\hat{\rho}(j))=\begin{cases}\alpha\hat{\rho}(j),&\text{if $n_{1}(j)/j>\hat{\rho}(j)$,}\\ \hat{\rho}(j),&\text{if $n_{1}(j)/j=\hat{\rho}(j)$,}\\ 1-\alpha(1-\hat{\rho}(j)),&\text{if $n_{1}(j)/j<\hat{\rho}(j)$,}\end{cases}$
where $\hat{\rho}(j)$ is an estimated optimal proportion and $n_{1}(j)$ the number of patients allocated to arm 1 after the allocation of $j$ patients. Following the recommendation of Burman (1996), the parameter $\alpha$ is typically chosen within the range of $0.4$ to $0.7$. We choose $\alpha=0.5$.
Both DBCD and ERADE effectively drive the design towards the theoretical optimal allocation. These targeting procedures enable faster convergence to the desired allocation proportions compared to simpler methods. For a more detailed discussion of these designs, we refer interested readers to Rosenberger and Lachin (2016).
2.3 Simulation Results
In this section, we present simulation results for the Wald test under three allocation rules: CR ($\rho_{CR}$), Neyman ($\rho_{N_{1}}$), and RSHIR ($\rho_{R_{1}}$), using ERADE as the targeting procedure. A significance level of $\alpha=0.05$ was used for the Wald test and will be used for all future tables and figures in this paper. The simulations in this section were conducted with a minimal burn-in period of $B=4$, two patients per arm allocated using permuted block randomization, and each scenario was simulated 10,000 times, ensuring a Monte Carlo error of less than 0.5% for both power and type-I error rate (Morris et al., 2019). We use a burn-in of 2 patients per arm consistently throughout the paper. There is no consensus in the literature on the optimal burn-in size for response-adaptive trials (Tang et al., 2025). We intentionally use the minimal burn-in (2 patients per arm) to investigate the maximal extent of type-I error rate inflation.
Table 1 presents the results for type-I error rate across the entire parametric space. Additionally, it reports results for a fixed $p_{0}=0.2$ ($p_{0}=0.7$) under an increasing (decreasing) sequence of treatment effects to illustrate differences in power.
Testing with $Z_{1}$
$p_{0}$ $p_{1}$ Type-I Error Rate or Power $n_{1}/n$ ENS
$\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$
0.1 0.1 5.0% 68.2% 68.1% 0.5 (0) 0.47 (0.1477) 0.46 (0.1471) 5 5 5
0.2 0.2 5.9% 82.2% 80.0% 0.5 (0) 0.46 (0.1570) 0.48 (0.1525) 10 10 10
0.3 0.3 6.3% 72.0% 66.8% 0.5 (0) 0.47 (0.1432) 0.48 (0.1336) 15 15 15
0.4 0.4 6.2% 64.7% 53.0% 0.5 (0) 0.47 (0.132) 0.48 (0.1063) 20 20 20
0.5 0.5 6.4% 61.9% 38.6% 0.5 (0) 0.47 (0.1261) 0.49 (0.0757) 25 25 25
0.6 0.6 6.0% 65.0% 26.6% 0.5 (0) 0.47 (0.1326) 0.49 (0.0490) 30 30 30
0.7 0.7 6.1% 71.9% 17.8% 0.5 (0) 0.47 (0.1438) 0.49 (0.0295) 35 35 35
0.8 0.8 6.1% 82.1% 10.6% 0.5 (0) 0.47 (0.1567) 0.49 (0.0127) 40 40 40
0.9 0.9 4.8% 68.3% 5.1% 0.5 (0) 0.47 (0.1477) 0.49 (0.0033) 45 45 45
[…]
0.2 0.3 15.3% 80% 76.7% 0.5 (0) 0.55 (0.1513) 0.57 (0.1420) 12.5 12.7 12.9
0.2 0.5 65.4% 88.3% 84.9% 0.5 (0) 0.62 (0.1342) 0.71 (0.0907) 17.5 19.3 20.7
0.2 0.7 97.1% 98.3% 97.8% 0.5 (0) 0.56 (0.1486) 0.79 (0.0476) 22.5 23.9 29.8
0.7 0.2 97.2% 98.5% 97.9% 0.5 (0) 0.38 (0.1408) 0.19 (0.0408) 22.5 25.5 30.2
0.7 0.4 62.1% 85.9% 72.0% 0.5 (0) 0.52 (0.1381) 0.33 (0.0507) 27.5 27.2 30.0
0.7 0.6 13.9% 70.4% 26.6% 0.5 (0) 0.51 (0.1394) 0.45 (0.0372) 32.5 32.5 32.8
0.7 0.8 15.0% 78.9% 22.0% 0.5 (0) 0.38 (0.1396) 0.52 (0.0213) 37.5 36.9 37.6
Table 1: Power or Type-I Error Rate, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for $n=50$, a burn-in size of $2$ patients per arm and different settings of $p_{0}$ and $p_{1}$.
We observe that both allocation strategies achieve their intended objectives. Using Neyman allocation results in the highest power, whereas RSHIR allocation maximizes the expected number of successes. However, it is important to note that the outcomes of these procedures exhibit high variability, particularly when Neyman allocation is applied.
As noted by Fleiss et al. (1981, page 30), the Wald test $Z_{1}$ can lead to type-I error rate inflation even when CR is used. Despite a Monte Carlo error of 0.5%, we observe type-I error rate inflation in both scenarios of $p_{0}=p_{1}=0.2$ and $p_{0}=p_{1}=0.7$, with an inflation exceeding 1% in the latter case. Furthermore, the use of optimal allocation exacerbates type-I error rate inflation to a level that is indeed surprising. Specifically, Neyman allocation $\rho_{N_{1}}$ inflates the type-I error rate to over 81% in the case of $p_{0}=p_{1}=0.2$, but the inflation with RSHIR allocation is almost equally high. The type-I error rate inflation across the whole parametric space is shown in Figure 1.
Figure 1: Comparison of Type-I error rates of: the Neyman proportion testing with the Wald test $\rho_{N_{1}}$, the RSHIR proportion testing with the Wald test $\rho_{N_{1}}$, complete randomization testing with the Wald test $\rho_{CR_{1}}$, complete randomization testing with the score test $\rho_{CR_{0}}$, the Neyman-like proportion testing with the score test $\rho^{n}_{N_{0}}$ and the RSHIR-like proportion testing with the score test $\rho^{n}_{R_{0}}$. In the both figures a burn-in size of $2$ patients per arm was used.
To the best of our knowledge, this issue has not been previously documented and renders the use of these procedures in their current form impractical.
3 Evaluating Existing Approaches to Reduce Type-I Error Rate Inflation
In this section, we present existing approaches that can help reduce the inflation of the type-I error rate. Although these methods have not been explicitly proposed for type-I error rate control, their application seems closely related, as they address issues that contribute to type-I error rate inflation, such as non-zero variance estimators or reduced adaptivity.
3.1 Agresti & Caffo Correction
Agresti and Caffo (2000) use the same test statistic $Z_{1}$ but modify the original MLE variance estimator by incorporating one artificial failure and one artificial success for each arm, adjusting the success probability estimates to $\hat{p}_{i}^{a}=(s_{i}+1)/(n_{i}+2)$, where $s_{i}$ is the number of successes. This adjustment ensures that the estimated success probabilities fall within the open interval $(0,1)$ and that therefore the estimated variances $\hat{p}_{i}^{a}(1-\hat{p}_{i}^{a})$ are non-zero.
However, the Agresti & Caffo estimator, $\hat{p}_{i}^{a}$, only converges to the MLE asymptotically. For small sample sizes, especially with success probabilities near 0 or 1, the Agresti & Caffo estimator $\hat{p}_{i}^{a}$ may introduce considerable bias and lead to incorrect test decisions. As shown in Table 2, while the correction reduces type-I error rate inflation, it does not control the type-I error rate at a significance level of $\alpha=0.05$.
$p_{0}$ $p_{1}$ Type-I Error Rate or Power ENS
Agresti & Caffo $B=12$ $B=30$ Agresti & Caffo $B=12$ $B=30$
$\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$
0.1 0.1 0.3% 0.1% 51.7% 52.5% 19.0% 19.0% 5 5 5 5 5 5
0.2 0.2 2.8% 0.4% 45.3% 44.3% 12.2% 12.1% 10 10 10 10 10 10
0.3 0.3 6.1% 1.3% 25.5% 24.5% 8.7% 9.0% 15 15 15 15 15 15
0.4 0.4 6.6% 3.0% 14.3% 14.3% 7.8% 7.7% 20 20 20 20 20 20
0.5 0.5 7.1% 5.3% 11.1% 9.3% 6.9% 7.1% 25 25 25 25 25 25
0.6 0.6 6.9% 7.5% 14.8% 6.7% 7.7% 6.6% 30 30 30 30 30 30
0.7 0.7 6.1% 11.5% 25.8% 6.0% 8.9% 6.2% 35 35 35 35 35 35
0.8 0.8 3.1% 9.1% 43.9% 5.7% 11.8% 5.4% 40 40 40 40 40 40
0.9 0.9 0.3% 3.4% 51.7% 4.8% 19.4% 4.4% 45 45 45 45 45 45
[…]
0.2 0.3 6.9% 2.3% 40.4% 41.1% 19.0% 20.1% 12.8 12.8 12.8 12.8 12.7 12.7
0.2 0.5 36.2% 25.1% 72.2% 71.3% 67.6% 67.6% 19.2 20.7 19.2 19.8 18.3 18.9
0.2 0.7 81.5% 75.8% 98.0% 97.4% 97.5% 97.4% 23.8 29.8 24.2 27.4 23.2 25.5
$p_{0}$ $p_{1}$ Type-I Error Rate or Power ENS
Agresti & Caffo $B=12$ $B=120$ Agresti & Caffo $B=12$ $B=120$
$\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$
0.1 0.1 0.6% 0.4% 76.5% 76.8% 7.9% 8.2% 20 20 20 20 20 20
0.2 0.2 3.7% 1.2% 44.8% 44.3% 5.9% 6.5% 40 40 40 40 40 40
0.3 0.3 8.4% 2.3% 23.2% 24.2% 5.7% 5.0% 60 60 60 60 60 60
0.4 0.4 7.0% 3.5% 12.9% 12.4% 5.3% 5.7% 80 80 80 80 80 80
0.5 0.5 7.1% 5.0% 9.7% 7.6% 5.5% 5.4% 100 100 100 100 100 100
0.6 0.6 7.2% 8.3% 13.6% 5.7% 5.7% 5.4% 120 120 120 120 120 120
0.7 0.7 8.0% 15.1% 24.4% 5.1% 5.5% 5.3% 140 140 140 140 140 140
0.8 0.8 3.7% 10.2% 44.2% 5.1% 6.1% 5.4% 160 160 160 160 160 160
0.9 0.9 0.8% 6.2% 76.2% 5.1% 7.7% 5.3% 180 180 180 180 180 180
[…]
0.2 0.3 14.1% 10.3% 56.3% 56.7% 38.5% 39.8% 51.2 51.9 51.8 51.9 50.6 51.0
0.2 0.5 49.6% 42.6% 99.5% 99.4% 99.5% 99.6% 78.1 84.3 78.4 81.5 73.3 76.4
0.2 0.7 89.7% 87.1% 100.0% 100.0% 100.0% 100.0% 97.0 121.9 98.6 112.7 93.3 102.7
Table 2: Comparison of type-I error rate or power and ENS for various settings of $p_{0}$ and $p_{1}$: The tables provide results for the Agresti & Caffo correction and different burn-in periods ($B=12$, $B=30$, $B=120$) for two sample sizes, $n=50$ (upper table) and $n=200$ (lower table). For the Agresti & Caffo correction a minimal burn of size $2$ patients per arm was used.
3.2 Ensuring Existence of Variance Estimators
Rosenberger et al. (2001) also suggest sampling with equal probability when one of the variance estimators is zero, indicating that all observations in that sample are identical. This often happens in small samples or at the beginning of a larger trial, especially with success rates near zero or one. Under these conditions, the adaptive design effectively simplifies to CR.
Testing with $Z_{1}$
$p_{0}$ $p_{1}$ Type-I Error Rate or Power $n_{1}/n$ ENS
$\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$
0.1 0.1 5.0% 4.8% 4.0% 0.5 (0) 0.49 (0.0036) 0.49 (0.0044) 5 5 5
0.2 0.2 5.9% 8.2% 8.1% 0.5 (0) 0.49 (0.0043) 0.49 (0.0062) 10 10 10
0.3 0.3 6.3% 8.6% 9.0% 0.5 (0) 0.49 (0.0023) 0.49 (0.0.005) 15 15 15
0.4 0.4 6.2% 7.2% 8.2% 0.5 (0) 0.49 (0.0008) 0.49 (0.0032) 20 20 20
0.5 0.5 6.4% 7.1% 7.4% 0.5 (0) 0.49 (0.0004) 0.49 (0.0020) 25 25 25
0.6 0.6 6.0% 7.7% 6.2% 0.5 (0) 0.49 (0.0007) 0.49 (0.0013) 30 30 30
0.7 0.7 6.1% 8.6% 6.1% 0.5 (0) 0.49 (0.0022) 0.49 (0.0008) 35 35 35
0.8 0.8 6.1% 7.9% 6.3% 0.5 (0) 0.49 (0.0042) 0.49 (0.0005) 40 40 40
0.9 0.9 4.8% 4.7% 5.0% 0.5 (0) 0.49 (0.0036) 0.49 (0.0003) 45 45 45
[…]
0.2 0.3 15.3% 18.1% 18.4% 0.5 (0) 0.53 (0.0032) 0.54 (0.0054) 12.5 12.6 12.7
0.2 0.5 65.4% 67.6% 68.5% 0.5 (0) 0.55 (0.0022) 0.61 (0.0036) 17.5 18.2 19.1
0.2 0.7 97.1% 97.3% 97.4% 0.5 (0) 0.53 (0.0032) 0.64 (0.0029) 22.5 23.2 26.0
Table 3: Results for sampling with equal probability if one of the two variance estimators is 0. Power or Type-I Error Rate, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for $n=50$, a burn-in size of $2$ patients per arm and different settings of $p_{0}$ and $p_{1}$.
Table 3 shows that adopting equal-probability sampling reduces type-I error rate inflation from 80% to 8% when success rates are not at extremes. However, this adjustment brings power and ENS metrics Neyman and RSHIR allocation closer to CR.
Hence, for smaller samples, this adjustment effectively attenuates type-I error rate inflation by reducing the adaptive design to CR and (even for larger samples). However, as shown in Table 3 it does not fully eliminate type-I error rate inflation and it considerably reduces any advantages from an adaptive approach.
3.3 Burn-In Periods to Stabilize Estimators and Decrease Adaptivity
A burn-in period is defined as the initial allocation of the first $B\in\{1,\dots,n\}$ patients with equal probability to each arm. To ensure an equal number of patients ($B/2$) in each arm, we enforce balance using methods such as simple random allocation or permuted block randomization. This step stabilizes the MLEs before adapting the allocation probabilities to reduce the likelihood of assigning patients in the wrong direction. The minimal burn-in size for RSHIR is one patient per arm ($B=2$), whereas for Neyman, it is two patients per arm ($B=4$), see Table 1. Table 2 demonstrates that (as the other approaches) the procedures with using larger burn-in periods still result in inflated type-I error rates. Moreover, increasing the burn-in size comes at the cost of diminishing patient-benefit gains (RSHIR) or power gains (Neyman). For Table 2 we used a burn-in period of six patients per arm ($B=12$), as recommend by Rosenberger et al. (2001), and 60% of the total trial size to demonstrate the the asymptotic behavior. In smaller trials, this trade-off often forces a near complete convergence to a CR design in an attempt to control the type-I error rate, though full control may not always be achievable, see Section 3.4. In larger trials, the same burn-in size allows for adaptive sampling while retaining error control, see Table 2.
3.4 Using the Score Test: Testing with a Pooled Variance Estimator
Eberhardt and Fligner (1977) and Fleiss et al. (1981, page 30) observed that the Wald-type test statistic $Z_{1}$ often exhibits an inflated type-I error rate in non-adaptive settings. To mitigate this issue, they propose testing the null hypothesis $H_{0}:p_{0}=p_{1}$ using the score test statistic (Agresti, 2002, p. 14), which is defined as
$Z_{0}=\frac{\hat{p}_{1}-\hat{p}_{0}}{\sqrt{\hat{p}\hat{q}\left(\frac{1}{n_{0}}+\frac{1}{n_{1}}\right)}},$
where $\hat{p}$ is the overall proportion of successes in the trial, calculated as the total number of successes divided by the trial size $n$, and $\hat{q}=1-\hat{p}$. Comparing the CR results from Table 1 with those from Table 4 and Figure 1 confirms that this test provides superior control over the type-I error rate.
The score test $Z_{0}$, is commonly used (Agresti, 2002) and it can be shown that the squared statistic, $Z_{0}^{2}$, is equivalent to the widely used Pearson chi-squared test (Scott, 2020, p. 119). In contrast, the variance of the test statistic $Z_{1}$ that was used by Tschuprow (1923); Robbins (1952); Rosenberger et al. (2001) incorporates the individual estimates of $p_{0}$ and $p_{1}$ in the estimator of the variance of the test statistic, enabling the construction of confidence intervals that align with the test decision. While using the Wald test typically achieves higher power compared to the score test, it does so at the expense of inflating the type-I error rate, even in non-adaptive settings (Fleiss et al., 1981, page 30).
Rosenberger et al. (2001) suggest using the allocation proportion $\rho_{N_{1}}$ with the $Z_{0}$ test. However, combining $\rho_{N_{1}}$ with the score test does not effectively control the type-I error rate or maximize power, as shown in Table 4. This is because $\rho_{N_{1}}$ is optimized for $Z_{1}$, not $Z_{0}$. Similarly, the RSHIR allocation proportion $\rho_{R_{1}}$, which is derived using variance of the Wald test, also fails with $Z_{0}$.
Testing with $Z_{0}$
$p_{0}$ $p_{1}$ Type-I Error Rate or Power $n_{1}/n$ ENS
$\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N^{n}_{0}}$ $\rho_{R^{n}_{0}}$ $\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N^{n}_{0}}$ $\rho_{R^{n}_{0}}$ $\rho_{CR}$ $\rho_{N_{1}}$ $\rho_{R_{1}}$ $\rho_{N^{n}_{0}}$ $\rho_{R^{n}_{0}}$
0.1 0.1 4.6% 0.3% 0.1% 6.8% 6.0% 0.5 (0) 0.47 (0.1477) 0.46 (0.1471) 0.49 (0.0047) 0.49 (0.0036) 5 5 5 5 5
0.2 0.2 5.5% 2.9% 0.4% 5.0% 5.2% 0.5 (0) 0.46 (0.1568) 0.48 (0.1525) 0.49 (0.0033) 0.48 (0.1525) 10 10 10 10 10
0.3 0.3 5.6% 6.7% 1.2% 4.9% 4.7% 0.5 (0) 0.47 (0.1432) 0.48 (0.1336) 0.49 (0.0015) 0.49 (0.0005) 15 15 15 15 15
0.4 0.4 6.2% 7.0% 2.9% 5.3% 4.8% 0.5 (0) 0.47 (0.132) 0.48 (0.1063) 0.49 (0.0006) 0.49 (0.0005) 20 20 20 20 20
0.5 0.5 6.0% 6.7% 5.4% 5.5% 4.6% 0.5 (0) 0.47 (0.1261) 0.49 (0.0757) 0.49 (0.0003) 0.49 (0.0017) 25 25 25 25 25
0.6 0.6 5.5% 7.2% 8.6% 5.3% 5.0% 0.5 (0) 0.47 (0.1326) 0.49 (0.0490) 0.49 (0.0006) 0.49 (0.0040) 30 30 30 30 30
0.7 0.7 5.6% 6.7% 13.2% 4.8% 5.5% 0.5 (0) 0.47 (0.1437) 0.49 (0.0293) 0.49 (0.0015) 0.49 (0.0074) 35 35 35 35 35
0.8 0.8 5.6% 3.2% 10.3% 5.3% 6.8% 0.5 (0) 0.47 (0.1567) 0.49 (0.0127) 0.49 (0.0033) 0.48 (0.0126) 40 40 40 40 40
0.9 0.9 5.1% 0.3% 5.5% 6.1% 7.9% 0.5 (0) 0.47 (0.1477) 0.49 (0.0033) 0.49 (0.0047) 0.47 (0.0145) 45 45 45 45 45
[…]
0.2 0.3 13.9% 7.0% 2.1% 12.5% 12.1% 0.5 (0) 0.56 (0.1488) 0.58 (0.1391) 0.46 (0.0023) 0.47 (0.0011) 12.6 12.8 12.9 12.3 12.4
0.2 0.5 64.1% 35.4% 24.7% 60.7% 60.0% 0.5 (0) 0.62 (0.1337) 0.71 (0.0913) 0.43 (0.0016) 0.50 (0.0020) 17.5 19.4 20.8 16.5 17.6
0.2 0.7 97.0% 84.2% 80.3% 96.4% 96.2% 0.5 (0) 0.55 (0.1494) 0.79 (0.0472) 0.46 (0.0023) 0.60 (0.0050) 22.5 23.8 29.8 21.5 25.1
Table 4: Comparison of five testing procedures using the score test: Type-I Error Rate or Power, proportion allocated to the treatment arm $n_{1}/n$, and expected number of successes (ENS) for $n=50$ and a burn-in size of $2$ per arm.
4 Revisiting Optimal RAR Using an Error Control Approach
Implementation of optimal allocation proportions requires estimating the unknown parameters $p_{0}$ and $p_{1}$. These parameters are estimated using maximum likelihood estimation. Using DBCD or ERADE the resulting procedure converges to the target theoretical optimal allocation $\rho$, e.g. $\rho_{N_{1}}$ or $\rho_{R_{1}}$.
The RAR procedures first introduced in this section substitutes the MLEs $\hat{p}_{0}$ and $\hat{p}_{1}$ for the unknown parameters $p_{0}$ and $p_{1}$ in the optimization problems from Section 2.1 directly, finding the optimal proportion at each step of the trial, which we denote as $\rho^{n}$ since the proportion depends on the sample size $n$. Hence, we take a finite sample approach based on the estimators rather than an asymptotic approach based on the unknown true values.
Moreover, we will use the variance of the score test $Z_{0}$ rather than $Z_{1}$. Since the test statistic of the score test $Z_{0}$ has an estimated variance that depends on the estimator $\hat{p}$ of unknown value of $p$ under the null hypothesis, we leverage this estimated variance as the basis for defining our optimization problem. This approach enables us to optimize the behavior of the score test for departures from the null hypothesis while controlling type-I error rate.
4.1 Neyman-like Allocation
In Section 3.4 we demonstrated that we cannot simply use $\rho_{N_{1}}$ as an allocation rule to maximize power and test with the score test $Z_{0}$ to control the type-I error rate. To address this limitation, we propose deriving a Neyman-like allocation tailored for the $Z_{0}$ test, minimizing its variance. First, we note that the estimator $\hat{p}$ is simply the total number of successes divided by the trial size $n$, but can be expressed using the weighted average of the MLEs of each treatment arm $\hat{p}:=(n_{0}\hat{p}_{0}+n_{1}\hat{p}_{1})/n$. We can then rewrite the squared variance of the test statistic as follows
$\frac{n\hat{p}\hat{q}}{n_{0}n_{1}}$ $=\frac{n\left(\frac{n_{0}\hat{p}_{0}+n_{1}\hat{p}_{1}}{n}\right)\left(1-\frac{n_{0}\hat{p}_{0}+n_{1}\hat{p}_{1}}{n}\right)}{n_{0}n_{1}}$
$=\frac{\hat{p}_{0}}{\rho^{n}n}+\frac{\hat{p}_{1}}{(1-\rho^{n})n}-\frac{(1-\rho^{n})\hat{p}_{0}^{2}}{\rho^{n}n}-\frac{2\hat{p}_{0}\hat{p}_{1}}{n}-\frac{\rho^{n}\hat{p}_{1}^{2}}{(1-\rho^{n})n}.$ (2)
We then derive the allocation proportion $\rho^{n}$ by minimizing the variance with respect to $\rho^{n}$ and solving for it, which yields
$\rho^{n}_{N_{0}}=\frac{\sqrt{\hat{p}_{0}\hat{q}_{0}}}{\sqrt{\hat{p}_{0}\hat{q}_{0}}+\sqrt{\hat{p}_{1}\hat{q}_{1}}}.$
Interestingly, it follows that $\rho^{n}_{N_{0}}=1-\hat{\rho}_{N_{1}}$ where $\hat{\rho}_{N_{1}}$ is the optimal solution form Section 2 with plugged-in estimators $\hat{p}_{0}$ and $\hat{p}_{1}$.
4.2 RSHIR-like Allocation
In this section we derive a proportion that minimizes the number of expected failures given a fixed variance constraint when testing with $Z_{0}$.
The second optimization problem from Section 2.1 remains unchanged but the variance in the constraint is now the variance of test statistic $Z_{0}$, specifically we can use it as formulated in equation (2) and rewrite the optimization problem to
$\text{min}_{\rho^{n}}\,\frac{\left((1-\rho^{n})(1-\hat{p}_{0})+\rho^{n}(1-\hat{p}_{1})\right)\left(\frac{\hat{p}_{0}}{\rho^{n}}+\frac{\hat{p}_{1}}{(1-\rho^{n})}-\frac{(1-\rho^{n})\hat{p}_{0}^{2}}{\rho^{n}}-2\hat{p}_{0}\hat{p}_{1}-\frac{\rho^{n}\hat{p}_{1}^{2}}{(1-\rho^{n})}\right)}{C}.$
deriving with respect to $\rho^{n}$ yields
$0=\frac{1}{C}\Bigg{[}(\hat{p}_{0}-\hat{p}_{1})\left(\frac{\hat{p}_{0}(1-\hat{p}_{0}+\rho^{n}\hat{p}_{0})}{\rho^{n}}+\frac{\hat{p}_{1}-\rho^{n}\hat{p}_{1}^{2}}{1-\rho^{n}}-2\hat{p}_{0}\hat{p}_{1}\right)$
$+(1-\hat{p}_{0}+\rho^{n}\hat{p}_{0}-\rho^{n}\hat{p}_{1})\left(\frac{\hat{p}_{1}(1-\hat{p}_{1})}{(1-\rho^{n})^{2}}-\frac{\hat{p}_{0}(1-\hat{p}_{0})}{(\rho^{n})^{2}}\right)\Bigg{]}.$
Even though we cannot find a close form solution for $\rho^{n}$ we can solve for it computationally and obtain $\rho^{n}_{R_{0}}$ (see Supplementary Material). In Figure 2 we demonstrate how the proportions differ theoretically and present simulation results in the next Section.
Figure 2: Optimal proportions $\rho^{n}_{R_{0}}$, $\rho_{R_{1}}$, $\rho^{n}_{N_{0}}$, and $\rho_{N_{1}}$ given $p_{0}=0.3,0.5,0.9$ and $p_{1}\in(0,1)$.
Figure 2 illustrates how the proportions change as $p_{1}$ varies, with $p_{0}$ fixed at 0.2. As indicated by the formulas, we observe that $\rho_{N_{0}}$ mirrors $\rho_{N_{1}}$. Notably, the curves for $\rho_{R_{0}}$ and $\rho_{R_{1}}$ suggest that patient allocation strategies aimed at optimizing patient outcomes differ significantly depending on the definition of the optimization problem. These differences imply that the regions where patient-benefit and statistical power can be improved vary with the choice of the test statistic as well. Specifically, when the curves for $\rho_{R_{0}}$ and $\rho_{N_{0}}$ are sufficiently close, there is potential to both increase power and allocate more patients to the superior treatment arm. It is important to note that these regions shift based on the value of $p_{0}$.
4.3 Simulation Results
Building on the theoretical results, we now analyze the simulation for the new allocation proportions. Figure 1 shows that the score test generally yields a lower type-I error rate than the Wald test, as seen in the comparison of CR for both tests ($\rho_{CR_{0}}$ vs. $\rho_{CR_{1}}$). The optimal optimal proportion $\rho^{n}_{N_{0}}$ successfully controls the type-I error rate and even outperform CR in this regard while the optimal proportion $\rho^{n}_{R_{0}}$ shows similar (or even better) performance for majority of the parametric space but is slightly inflated for values around $0.8$.
Table 4 shows that $\rho^{n}_{N_{0}}$ and $\rho^{n}_{R_{0}}$ achieve similar power to CR, with differences within few percent in either direction across parameter combinations. The difference in type-I error rate control varies. The Neyman-like optimal allocation $\rho^{n}_{N_{0}}$ generally provides better control of the type-I error rate compared to CR, except for values near the boundaries (i.e., close to 0 or 1). The RSHIR-like allocation proportion $\rho^{n}_{R_{0}}$ shows even better type-I error rate control for the majority of the parametric space, but the type-I error rate is inflated earlier for values approaching 1. While $\rho^{n}_{N_{0}}$ may reduce the ENS compared to CR, $\rho^{n}_{R_{0}}$ tends to increase it and typically allocates more patients to the superior arm.
5 Guidance from Redesigning Trial Examples
5.1 Early Phase Trial
We redesign a study to evaluate the statistical properties of an early-phase, two-armed trial designed to assess the efficacy of a treatment compared to a control. Our study was based on survival outcomes observed in a multicenter randomized phase 2 trial investigating the impact of neoadjuvant chemotherapy (NAC) in patients with locally advanced esophageal cancer (EC). The trial compared patients who received two courses of NAC versus those who received three courses, revealing that a tumor reduction rate lower than 10% during the third course was associated with unfavorable overall survival (OS). Specifically, we use the 2-year OS rate of 89.3% in the more favorable group and 63.5% in the less favorable group as the basis for our assumed treatment and control arm success rates, respectively Kubo et al. (2023). The trial was designed with a total sample size of 68 patients.
Results are summarized in Table 5. Among the allocation proportions examined, the RSHIR-like allocation $\rho^{n}_{R_{0}}$, the Neyman-like allocation $\rho^{n}_{N_{0}}$, and complete randomization, when testing with $Z_{0}$, successfully control the type-I error rate. Complete randomization when testing with the Wald test exhibited slight inflation of type-I error. However, the original Neyman $\rho_{N_{1}}$ and RSHIR $\rho_{R_{1}}$ allocations resulted in highly inflated type-I error rates, making them impractical for use.
Our findings suggest that both the RSHIR-like and Neyman-like allocation are viable methods, as they properly control the type-I error rate. While they have slightly lower power (less thatn $1\%$ difference) than complete randomization, they provide a significant increase in the expected number of successes, making them preferable approaches in this context.
Test Procedure Type-I error Power % sup Arm (Var) ENS
NAC: $N=68$, $p_{0}=0.635$, $p_{1}=0.893$
$Z_{1}$ $\rho_{CR}$ 5.5% 75.8% 0.5 (0) 51.9
$Z_{0}$ $\rho_{CR}$ 4.7% 74.2% 0.5 (0) 51.9
$Z_{1}$ $\rho_{N_{1}}$ 65.7% 94.2% 0.2216 (0.1024) 47.1
$Z_{1}$ $\rho_{R_{1}}$ 23.0% 76.6% 0.5798 (0.0207) 53.4
$Z_{0}$ $\rho^{n}_{N_{0}}$ 4.6% 73.6% 0.6064 (0.0033) 53.8
$Z_{0}$ $\rho^{n}_{R_{0}}$ 4.9% 73.4% 0.6909 (0.0076) 55.3
CALISTO: $N=1502$, $p_{0}=0.941$, $p_{1}=0.991$
$Z_{1}$ $\rho_{CR}$ 5% 100% 0.5 (0) 1450.9
$Z_{0}$ $\rho_{CR}$ 5% 100% 0.5 (0) 1450.9
$Z_{1}$ $\rho_{N_{1}}$ 96.3% 99.9% 0.1329 (0.1042) 1423.4
$Z_{1}$ $\rho_{R_{1}}$ 5.4% 100% 0.5073 (0.0005) 1451.5
$Z_{0}$ $\rho^{n}_{N_{0}}$ 5.2% 100% 0.7139 (0.0014) 1467
$Z_{0}$ $\rho^{n}_{R_{0}}$ 5.1% 100% 0.8298 (0.0031) 1475.7
Table 5: Results for two different sample sizes. The first set corresponds to $N=68$, $p_{0}=0.635$, $p_{1}=0.893$ with a number of simulations $10^{4}$ and minimal burn-in of $2$ patients per arm. The second set corresponds to $N=1502$, $p_{0}=0.941$, $p_{1}=0.991$ with the same simulation settings.
5.2 Confirmatory Trial: CALISTO
In this section, we redesign a published clinical trial to demonstrate the effect of the newly proposed proportions. We examine the CALISTO trial (Decousus et al., 2010). The goal of this trial was to assess the efficacy of a new drug, Arixtra, compared to a placebo in patients with acute symptomatic lower limb thrombophlebitis. Success was defined as the absence of a composite set of events, including death from any cause, symptomatic pulmonary embolism, symptomatic deep-vein thrombosis, symptomatic extension to the saphenofemoral junction, and symptomatic recurrence of superficial-vein thrombosis by day 47.
The trial results revealed a success rate of 94.1% in the placebo group and 99.1% in the treatment group. Originally, the study was designed to include 3002 patients, but only $n=1502$ patients were included in the analysis. We compare optimal proportions aimed at minimizing failures $\rho_{R_{1}}$ and $\rho^{n}_{R_{0}}$, as well as maximizing power, $\rho_{N_{1}}$ and $\rho^{n}_{N_{0}}$ to CR. For each method we use the test statistic $Z_{0}$ or $Z_{1}$ that was used in the optimization problem but for CR we present both. For each method, a burn-in period of $2$ patients per arm was implemented.
Table 5 demonstrates that CR does lead to asymptotic type-I error control using either the Wald or score test. The proportions $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ almost control type-I error rate, while $\rho_{R_{1}}$ and more so $\rho_{N_{1}}$ do not. While all designs have similar power, both proportions $\rho^{n}_{R_{0}}$ and $\rho^{n}_{N_{0}}$ are able to increase patient benefit significantly. Especially, using $\rho^{n}_{R_{0}}$ leads to 25 additional success compared to CR.
6 Discussion and Future Research
In this paper, we introduce two optimal allocation proportions that demonstrate type-I error rate control in adaptive two-armed trials with binary outcomes. In finite samples they achieve type-I error rate control for the largest margin of the parametric space (90%). These proportions offer a promising balance between statistical rigor and practical relevance. Our proposed methods demonstrate power performance that is comparable to existing approaches, with occasional gains, particularly in small sample settings. Importantly, the proposed proportions have the advantage of increasing the likelihood that patients receive the superior treatment.
A potential concern with using estimated quantities rather than true values in the optimization process could be the risk of introducing additional variability. However, as we show in Section 4, this concern is unfounded; there is no meaningful increase in variability when using estimators in our proposed framework. This finding strengthens the practical applicability of our method by demonstrating that it maintains robust statistical properties even when applied in real-world scenarios where true parameters are unknown.
Despite the promising results, some limitations remain. We focused on two-armed trials with binary endpoints, as these are relevant in practice (Pin et al., 2025) and are among the most frequently discussed settings in the theoretical literature (Hu and Rosenberger, 2006). Azriel et al. (2012) discusses the relationship to other optimal allocations for discrete and continuous endpoints that optimize for power. However, the scope of our results do not yet extend to multi-armed trials, which remain an important area for future research.
While our focus in this paper is on two-arm trials, there is a large literature on optimal allocation proportions in multi-armed settings. However, the multi-armed case poses additional challenges (Tymofyeyev et al., 2007; Sverdlov et al., 2020). For instance, multi-armed extensions of Neyman allocation exist (Tymofyeyev et al., 2007), though they tend to allocate patients only to the best and worst arms, which is undesirable in practice. As shown in Zhu and Hu (2009); Baldi Antognini et al. (2021), additional constraints (e.g., minimum allocation proportions) are needed to ensure all arms are represented. Other authors have proposed allocations tailored to specific hypothesis tests or altered objectives (Jeon and Hu, 2010; Azriel and Feigin, 2014; Biswas et al., 2011). Our proposed proportions, which are optimized specifically for the score test in the two-arm setting, likely extend in the same way the multi-arm case but also have the same limitations. While a score-test analog of the multi-armed Neyman allocation could, in principle, be derived, it would also only allocate to two arms unless further constraints are imposed. Tymofyeyev et al. (2007) investigated multi-armed extensions of both Neyman and RSHIR allocation, but the latter solution requires the choice of a smoothing kernel function and the choice of unknown parameters, but do not provide a closed-form solution.
An alternative route of approaching type-I error rate control for optimal proportions would be to incorporate it directly as a constraint in the optimization problem. This approach has been discussed, for example in Baas et al. (2024), but this would require separate computation for each sample size $n$ over the entire parameter space. Instead, we aim to highlight the practical benefits of using better suited test statistics within standard adaptive design frameworks.
Lastly, binary endpoints, such as success or failure, are a special case in that their variance depends directly on the mean, i.e., the success rate. This influences the behavior of optimal allocation rules and differentiates binary trials from most trials with continuous outcomes. While our work addresses this specific case, optimal proportions for other endpoints will look different and do not have the same properties. Ye et al. (2024) explored the robustness of RAR and the use of the Wald test for continuous responses and large samples.
There are several promising directions for future work. First, we want to adapt our approach to alternative measures of interest, such as relative risk or odds ratios. The calculations are straightforward for the Wald test (Rosenberger et al., 2001; Pin et al., 2024) and can be equivalently derived for the score test. However, these measures are widely used in practice and present different statistical properties that may influence optimal allocation strategies. Second, we aim to explore optimal allocation proportions for nonparametric tests and exact tests, which are particularly relevant in small-sample trials where traditional asymptotic methods may not perform well. Additionally, one could investigate deriving optimal proportions for finite sample corrections, such as the Agresti and Caffo correction, i.e. using the adjusted estimators in the optimization problem.
This work represents a significant step towards broader implementation of these methods. It addresses the current regulatory landscape, which prioritizes both patient-centered trials and robust type-I error rate control, by providing a framework that achieves both.
Acknowledgments
The authors acknowledge the use of large language models (LLMs) to assist with generating figures and refining grammar and wording in this paper. The LLMs were not used for data analysis, interpretation, or original scientific writing. All content has been carefully reviewed and verified by the authors, who take full responsibility for the integrity and accuracy of the work.
The authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP) and the Cusanuswerk e.V. (LP). SSV is part of PhaseV’s advisory board. Additionally, LP’s visit to the Department of Statistics at George Mason University to collaborate with WR on this project was supported by the Trials Methodology Research Partnership Doctoral Training Program and the Fisher Memorial Trust. LP is thankful for the departments hospitality. We thank Dr. Stef Baas for revising our codes and spotting two minor errors. The authors are grateful to the AE and referees for their insightful suggestions.
Supplementary Material (Code)
The code referenced in Section 4.2 is available with this paper at the Biometrics website on Oxford Academic. The code can be found at the GitHub repository https://github.com/lukaspinpin/Optimal-Proportions, and contains the R code necessary to replicate the results presented in the paper.
Data Availability
There are no data presented in this paper.